A car goes from rest to 50 m/s in 20 s, what is its AVERAGE ACCELERATION.

b) What is that ACCELERATION as a fraction of g (acceleration of gravity)?
Answer #1

Given,

The initial velocity, vi = 0 m/s (rest)

The final velocity, vf = 50 m/s

The time, t = 20 sec

The acceleration, a = v / t

a = (vf - vi) / t

= (50 - 0) / 20

= 2.5 m/s2